# Interpretable Visualizations of Data Spaces for Classification Problems

### v3

@@ III.3 Understanding feature importance in non-linear classification of metals/nonmetals
In general, the SHAP values and random forest feature importances correlate strongly with the representation of the features in the first principal covariate. However, interpreting models using these techniques has limitations. First, for these feature importance metrics to be useful, it is necessary that the features themselves can be interpreted. While this is not the case here, when abstract descriptors are used, such as the RDKit fingerprint in it will be difficult to relate these metrics to chemical concepts, which will make QSPR unachievable. With PCovC, model interpretation does not have to be limited to features the model uses for training, as visualizing or computing correlation coefficients can be performed trivially with the PCovC latent space. In addition, it is difficult to determine the span of the feature’s importance across the space of observations using solely feature importance metrics. For example, consider the map showing the average number of p electrons in the KPCovC projection in 8. This map shows that there is a strong population of metals with a small number of p electrons, but this population does not span the entire population of metals. While SHAP can similarly achieve sample-level resolution in this manner, PCovC can also be used to identify subgroups across the decision boundary, such as the sets of materials with high and low electronegativities in Fig. 8.
@@ III.4 Enhancing nonlinear dimensionality reduction techniques with MNIST
Nonlinear methods are often essential when facing problems in which local correlations between features dictate class labels. Consequently, with the MNIST handwritten digits dataset, dimensionality reduction algorithms that optimize the preservation of local neighborhoods of points in their reduced spaces, such as t-SNE and UMAP, can find embeddings with far cleaner digit separation than PCA or even LDA [7]. This advantage, however, is balanced by an increased computational cost. For this reason, when working with large datasets, many have considered using PCA as a preprocessing step to reduce data and make nonlinear dimensionality reduction feasible [47, 48]. At the same time, linear classification models can achieve a respectable performance with MNIST; despite its simplicity, a logistic regression model trained on the flattened images can achieve an accuracy of 92%. This begs the question: if there exists linearly decodable information that can be used to partially separate classes, can this information be used to enhance a nonlinear mapping strategy?
Figure 9: PCovC and PCA-reduced UMAP projections on the MNIST dataset. (top) Accuracy of a logistic regression model trained in PCA, PCovC, PCovC-reduced UMAP, and PCA-reduced UMAP latent spaces. (middle row) UMAP projections of MNIST obtained from PCA-reduced latent spaces. (bottom row) UMAP projections of MNIST obtained from PCovC-reduced latent spaces. In the titles, we indicate the accuracy of a logistic regression model on the resulting 2D map for the testing data. Color corresponds to the digit label and the backgrounds show the estimated decision boundaries.
In Fig. 9, we explore the use of PCovC as a replacement to PCA as a preprocessing step to UMAP. We observe that UMAP projections learned from PCovC-reduced subspaces of the data result in clearer digit separation than those learned from PCA-reduced subspaces. In practice, this means that PCovC can be used in two ways: PCovC can substitute PCA at a fixed subspace dimensionality to enhance the quality of the subsequent UMAP embedding without a significant increase in the computational cost, or it can be used to achieve stronger dimensionality reduction prior to UMAP without sacrificing class delineation in the resulting projection.
@@ IV Conclusions
In this work, we demonstrated that hybrid supervised-unsupervised mapping techniques, namely principal covariates classification, can be transformative tools for visualizing data spaces through the lens of classification problems. We observed that, compared to PCA and LDA, combining classification performance and information preservation in a single mapping strategy can unveil decision boundaries while still preserving similarity between data points, and can even lead to improvements in classification performance. In computational chemical analysis, these aspects can lead to furthering our ability to identify and interpret structure-property relationships.
@@ A.1 Principal Covariates Classification
Behind the curtain, obtaining a PCovC map requires a few key steps. First, we must obtain $\mathbf{Z}$ for use in Eq. (2) through fitting a linear classifier with $\mathbf{X}$ and $\mathbf{Y}$. The exact choice of linear classification model is flexible. The projector from the feature space to the latent space can be mathematically derived through following the procedure in [33], but with the evidence matrix $\mathbf{Z}$ used instead of the linear regression approximation $\mathbf{\hat{Y}}$. When $n_{\mathbf{features}}>n_{\mathbf{samples}}$, one should obtain the projectors by performing an eigendecomposition of a modified Gram matrix $\mathbf{\tilde{K}}$ (Eq.  3).
@@
Selecting $\alpha=1$ will still result in a standard PCA decomposition; selecting $\alpha=0$, on the other hand, does not have any intuitive analog, unless if using linear regression to directly predict class likelihoods. In that case, we have $\mathbf{Z}=\mathbf{\hat{Y}}$ and $\mathbf{P}_{TZ}$ will be the linear regression solution of $\mathbf{T}$ onto the output $\mathbf{Y}$, thus, PCovC at $\alpha=0$ will produce the same result as a linear regression.

### v4

@@ III.3 Understanding feature importance in nonlinear classification of metals/nonmetals
In general, the SHAP values and random forest feature importances correlate strongly with the representation of the features in the first principal covariate. However, interpreting models using these techniques has limitations. First, for these feature importance metrics to be useful, it is necessary that the features themselves can be interpreted. While this is not the case here, when abstract descriptors are used, such as the RDKit fingerprint in Sec. III.1, it will be difficult to relate these metrics to chemical concepts, which will make QSPR unachievable. With PCovC, model interpretation does not have to be limited to features the model uses for training, as visualizing or computing correlation coefficients can be performed trivially with the PCovC latent space. In addition, it is difficult to determine the span of the feature’s importance across the space of observations using solely feature importance metrics. For example, consider the map showing the average number of p electrons in the KPCovC projection in 8. This map shows that there is a strong population of metals with a small number of p electrons, but this population does not span the entire population of metals. While SHAP can similarly achieve sample-level resolution in this manner, PCovC can also be used to identify subgroups across the decision boundary, such as the sets of materials with high and low electronegativities in Fig. 8.
@@ III.4 Enhancing nonlinear dimensionality reduction techniques with MNIST
Figure 9: Comparison between PCA, PCovC, and UMAP on the MNIST dataset. (top row) Maps according to the first and second PCA, PCovC, and UMAP dimensions. In the lower left, we denote the accuracy of a logistic regression model on the resulting 2D map for the offset testing data. Color corresponds to the digit label, and opacity corresponds to the training (translucent) versus testing (opaque) split.
Nonlinear methods are often essential when facing problems in which local correlations between features dictate class labels. Consequently, with the MNIST handwritten digits dataset, dimensionality reduction algorithms that optimize the preservation of local neighborhoods of points in their reduced spaces, such as t-SNE and UMAP, can find embeddings with far cleaner digit separation than linear models [33]. With MNIST, a logistic regression model trained on the flattened images can achieve an accuracy of 92%. However, as can be seen in Fig. 9, UMAP significantly outperforms PCovC in a latent space linear classification task, despite the fact that UMAP is both unsupervised and does not enforce linear cluster separation. This failure occurs because, while linear classification is possible in the high-dimensional space, the locally defined class structure requires more than two dimensions to be resolved by PCovC’s linear projection. This advantage, however, is balanced by an increased computational cost. For this reason, when working with large datasets, many have considered using PCA as a preprocessing step to reduce data and make nonlinear dimensionality reduction feasible [57, 11]. This begs the question: if there exists linearly decodable information that can be used to partially separate classes, can this information be used to enhance a nonlinear mapping strategy?
Figure 10: UMAP projections on the MNIST dataset with PCovC and PCA pre-processing. (top) Accuracy of a logistic regression model trained in PCA, PCovC, 2D UMAP on PCovC components, and 2D UMAP on PCA components. (middle row) UMAP projections of MNIST obtained from PCA-reduced latent spaces. (bottom row) UMAP projections of MNIST obtained from PCovC-reduced latent spaces. In the titles, we indicate the accuracy of a logistic regression model on the resulting 2D map for the testing data. Color corresponds to the digit label and the backgrounds show the estimated decision boundaries.
In Fig. 10, we explore the use of PCovC as a replacement to PCA as a preprocessing step to UMAP. We observe that UMAP projections learned from PCovC-reduced subspaces of the data result in clearer digit separation than those learned from both the PCA-reduced subspaces and UMAP on the high-dimensional space. In practice, this means that PCovC can be used in two ways: PCovC can substitute PCA at a fixed subspace dimensionality to enhance the quality of the subsequent UMAP embedding without a significant increase in the computational cost, or it can be used to achieve stronger dimensionality reduction prior to UMAP without sacrificing class delineation in the resulting projection.
@@ IV Conclusions
In this work, we demonstrated that hybrid supervised-unsupervised mapping techniques, namely principal covariates classification, can be transformative tools for visualizing data spaces through the lens of classification problems. The corresponding open-source code is available at https://github.com/cersonsky-lab/PCovC. We observed that, compared to PCA and LDA, combining classification performance and information preservation in a single mapping strategy can unveil decision boundaries while still preserving similarity between data points, and can even lead to improvements in classification performance. In computational chemical analysis, these aspects can lead to furthering our ability to identify and interpret structure-property relationships.
@@ A.1 Principal Covariates Classification
Behind the curtain, obtaining a PCovC map requires a few key steps. First, we must obtain $\mathbf{Z}$ for use in Eq. (3) through fitting a linear classifier with $\mathbf{X}$ and $\mathbf{Y}$. The exact choice of linear classification model is flexible. The projector from the feature space to the latent space can be mathematically derived through following the procedure in [21], but with the evidence matrix $\mathbf{Z}$ used instead of the linear regression approximation $\mathbf{\hat{Y}}$. When $n_{\mathbf{features}}>n_{\mathbf{samples}}$, one should obtain the projectors by performing an eigendecomposition of a modified Gram matrix $\mathbf{\tilde{K}}$ (Eq.  4).
@@
Selecting $\alpha=1$ will still result in a standard PCA decomposition; selecting $\alpha=0$, on the other hand, can result in re-learning $\mathbf{P}_{TZ}$ when the dimensionality of $\mathbf{T}$ is the same as the dimensionality of $\mathbf{Z}$.
@@ B.3 SHAP Values and Saliency Maps
Appendix C Computational Performance Comparisons
To further evaluate the performance of PCovC, we ran a number of PCA, LDA, PCovC, t-SNE, and UMAP fits on synthetic data. PCA and LDA fits were performed using scikit-learn [36] , t-SNE fits were performed using openTSNE [38] with default hyperparameters, and UMAP fits were performed using umap-learn [33] with default hyperparameters. All benchmarks were ran on a server with AMD EPYC 7763 processors with 10 GB of RAM.
(a) Performance of PCovC, logistic regression, LDA, PCA, t-SNE, UMAP, and supervised UMAP with $n_{\mathbf{samples}}=1,000,10,000,$ and $100,000$. Here, we apply each method on a synthetic dataset with $n_{\mathbf{features}}=100$ and $n_{\mathbf{classes}}=2$.
(b) Performance of PCovC, logistic regression, LDA, PCA, t-SNE, UMAP, and supervised UMAP with $n_{\mathbf{features}}=50,100,1,000,$ and $5,000$. Here, we apply each method on a synthetic dataset with $n_{\mathbf{samples}}=10,000$ and $n_{\mathbf{classes}}=2$.
(c) Performance of PCovC, logistic regression, LDA, PCA, t-SNE, UMAP, and supervised UMAP with $n_{\mathbf{classes}}=2,10,50,$ and $100$. Here, we apply each method on a synthetic dataset with $n_{\mathbf{samples}}=10,000$ and $n_{\mathbf{features}}=2$.
Figure 11: Performance benchmarks of PCovC with existing methods.
Generally, these results indicate that PCovC performs competitively with existing linear dimensionality reduction algorithms and can be faster than UMAP or t-SNE, as long as the data is not extremely high-dimensional.
Appendix D Comparison with Existing Methods
Here, we compare the maps obtained with PCovC to other popular data visualization techniques. Maps were made using the data discussed in Sec. III.1, with the “tox21-ache-p3 (-1)" assay results as labels.
Figure 12: Comparison of PCovC to PCA, supervised PCA, t-SNE, unsupervised UMAP, and supervised UMAP for neurotoxicity classification. White and blue points indicate “non-toxic” and “toxic” molecules, respectively. Opacity denotes train (translucent)/ test (opaque) split. In each panel is a confusion matrix showing the accuracy of logistic regression on the testing data as they appear in the resulting map, where “TN” indicates the number of “true negatives”, “FP” indicates the number of “false positives”, and so on.